Circular No. 8805

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 C1 (CHRISTENSEN) E. J. Christensen reports his discovery of a comet on CCD images taken

with the Catalina Sky Survey’s 0.68-m Schmidt telescope (discovery observation tabulated below), the object displaying a 12′′–15′′ coma, slightly elongated toward p.a. 240^{o} on four co-added 30-s exposures obtained in average seeing and light cirrus. Christensen adds that four co-added 60-s exposures on Feb. 10.2 UT in average seeing show a 10′′ coma and 15′′ tail in p.a. 250^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, G. Sostero and E. Guido write that thirty co-added 180-s unfiltered CCD exposures taken on Feb. 10.05–10.09 with a 0.45-m f/4.4 reflector at Remanzacco, Italy, show a compact coma 12′′ in diameter and a short tail 16′′ long toward p.a. 260^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Feb. 9.28032  10^{h}16^{m}07.^{s}17  +27^{o}03′13′′.4  18.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-C50. It is possible that this comet is of short period.

T = 2007 June 3.159 TT  ω = 185.^{o}355
Ω = 27.726 }2000.0
q = 1.42180 AU  i = 9.092

COMETS C/2006 U14, C/2006 U15, C/2006 U16 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8788). K. Battams notes that C/2006 U14 was stellar in appearance and of mag ∼ 7 in C3 images, while it appeared “condensed” and small in C2 images. C/2006 U15 was very small, faint, and a little diffuse. C/2006 U16 (which was independently reported by T. Chen, H. Su, and R. Kracht) was stellar in appearance and very faint (mag ∼ 8).

Comet  2006 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2006 U14 Oct. 28.096 13^{h}57.^{m}5 −14^{o}42′ C3/2 HS 2007-B37
C/2006 U15  28.329 14 04.5 −14 15 C2  LC 2007-B37
C/2006 U16  30.604 14 13.9 −15 05 C2  TC 2007-B37
2007 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green